Use integration by parts to integrate ∫x secx tanx dx
1
Expert's answer
2020-03-09T10:56:24-0400

Solution:


"\int {x\sec x \tan x}~dx~~~\small{=>} ~~~|\sec x = \frac{1}{\small\cos x}|~~~ \small{=>}~~~\large\int \frac{\small{x\cdot\sin x}} {\small{\cos^2x}}\small{dx} ~~~~~~~~~~~~~~(1)"


After getting expression (1) we can use method integrate by parts: "\int v~du=u\cdot v - \int u~dv"

"\large\int \frac{\small{x\cdot\sin x}} {\small{\cos^2x}}\small{dx} =\begin{bmatrix}
 v = x, ~dv = dx~~~~~~~~~~~~~~~~~~\\
du=\frac{\small{\sin x}} {\small{\cos^2x}}dx, u^* = \frac{\small{1}} {\small{\cos x}}
\end{bmatrix} = \frac{\small{x}} {\small{\cos x}}-\int\frac{\small{dx}} {\small{\cos x}} = \frac{\small{ x}} {\small{\cos x}}-(**) + C;"


"(**)~~~\int\frac{\small{d x}} {\small{\cos x}} = \int \frac{\small{\cos x}} {\small{\cos^2x}}dx = 
\int \frac{\small{d(\sin x})} {\small{\cos^2x}} = \int \frac{\small{d(\sin x)}} {\small{1-\sin^2x}}~~\small{=>}~~|\sin x =t~|~~\small{=>}~~\int \frac{\small{dt}} {\small{1-t^2}}"


So, we get a tabular integrale: "\int \frac{\small{dt}} {\small{1-t^2}}=-\frac{1}{2}\ln|\frac{t-1}{t+1}|+C;"


"(**)~~~\int\frac{\small{d x}} {\small{\cos x}} =-\frac{1}{2}\ln\large{|}\frac{\small{\sin x -1}}{\small{\sin x + 1}}\large{|};"


"\large\int \frac{\small{x\cdot\sin x}} {\small{\cos^2x}}\small{dx} =\frac{\small{ x}} {\small{\cos x}}+\frac{1}{2}\ln\large{|}\frac{\small{\sin x -1}}{\small{\sin x + 1}}\large{|} + C;"


"u^*=\int \frac{\small{\sin x}} {\small{\cos^2x}}dx = -\int \frac{\small{d(\cos x)}} {\small{\cos^2x}}=\frac{\small{1}} {\small{\cos x}};"


Answer: "\large\int \frac{\small{x\cdot\sin x}} {\small{\cos^2x}}\small{dx} =\frac{\small{ x}} {\small{\cos x}}+\frac{1}{2}\ln\large{|}\frac{\small{\sin x -1}}{\small{\sin x + 1}}\large{|} + C;"